Use a standard normal distribution table to obtain the z-score that has an area of 0.995 to its right Round to two decimal places as needed.)Please check these thank you
Answer #1

a)

We have to calculate z sucth that

P( Z > z) = 0.995

That is

P( Z < z) = 0.005

From the Z table, z-score for the probability of 0.005 is -2.58

z = -2.58

b)

We have to calculate z such that

P( -z < Z < z) = 0.52

P( Z < z) - P( Z < -z) = 0.52

P( Z < z) - ( 1 - P( Z < z) ) = 0.52

P( Z < z) - 1 + P( Z < z) = 0.52

2 * P( Z < z) = 1.52

P( Z < z) = 0.76

From the Z table, z-score for the probability of 0.796 is 0.7063

The z-scores are -0.71 , 0.71 (Rounded to two deciaml palces)